OpenStudy (anonymous):

find the roots of the polynomial equation 2x^3+2x^2-19x+20=0

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

what are also the real and complex numbers

OpenStudy (anonymous):

http://www.mathsisfun.com/numbers/real-numbers.html

OpenStudy (anonymous):

but what are they for this equation

OpenStudy (anonymous):

the roots are as follow, x= -4, x = (6+\[\sqrt{-4}\])/4, x = (6-\[\sqrt{-4}\])/4

OpenStudy (anonymous):

as you cant take sq root of minus thereofre we instead of sqrt of -4 will be written as 2i where i is sqrt of -1

OpenStudy (anonymous):

you can write the roots x = -4, x=( 6+2i)/4, x=( 6-2i)/4

OpenStudy (anonymous):

the first root is a real number whereas the other two are complex